# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 3.3 G-emergence Theory
G-emergence theory proposed by Seth in 2008 Seth (2008) is one of the earliest works on quantitative measure of emergence. His basic idea is to use nonlinear granger causality to measure weak emergence in complex systems.
@@
Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$ along, then $B$ is said to be "Granger-cause" of $A$. This implies the existence of Granger causality between $A$ and $B$.
@@
When applying a bivariate autoregressive model for predictions, residual terms are included in the equations of the two variables. Then, the residuals can be utilized to quantify the extent of the causal effect in G-causality. The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals, one is residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other one is the residual of the full prediction model, which shown in Figure 3. Besides, in Seth (2008), the author also defined G-autonomous as if the past values of a time series can help to predict its own future values. And the degree of G-autonomous can be measured in a similar way as G-causality.
Figure 3: G-causality and G-autonomous
@@
With these two basic notions in G-causality, the author defines G-emergence on macro-variables in Figure 4: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if: (i) $M$ is G-autonomous with respect to $\{m\}$ and (ii) $\{m\}$ is G-cause of $M$. The degree of G-emergence can also be quantitatively measured by multiplying the degree of G-autonomous of $M$ and the average of G-causes of $\{m\}$ on $M$.
Figure 4: G-autonomous
@@
The author tests G-emergence theory on the Boid model. This model is a famous artificial life model to simulate the flocking behavior of birds by using three simple rules: cohesion, alignment, and separation Reynolds (1987). All these basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled. The author found that with the strength of cohesion increased, the G-emergence of the entire flock is also increased. Here, the macro-variable is selected as the center of the flock (center-of-mass), and each bird is a micro-variable. The author also discovered a downward causality phenomenon in this simple model that is the center of mass can be used to predict each individual bird. However, the author does not distinguish downward causation from other common causality in this work because Granger’s causality is not a real causal relationship.
@@
Seth’s G-emergence theory is the first attempt to quantify emergence phenomenon via causality measure. However, the causality measure that the author used is granger causality which is not a strict causal measure and it is also depends on the regression method to be used. Furthermore, the measure is defined on variables but not dynamics which means the result depends on the selection of variables.
@@ 3.4 Other Quantitative Theories of Emergence
There are alternative means of measuring emergence that do not require relying on causality. Two different ways are discussed by the scholars, one is to understand emergence as a process from disorder to order, another one is to understand emergence from the view of "the whole is greater than the sum of parts".
For example, Moez Mnif and Christian Müller-Schloer used Shannon’s entropy  Mnif and Müller-Schloer (2011) to measure order and disorder. In a self-organized process, emergence occurs when there is an increase in order. This increase can be quantified by measuring the Shannon entropy difference between the initial state and the final state, denoted as $H_{start}-H_{end}$. However, this definition has two limitations. Firstly, the measurement of entropy, denoted as $H$, is dependent on the abstract level of observation. Therefore, it is necessary to account for the entropy increase resulting from a change in the observer’s abstract level. Secondly, the choice of the initial condition of the system is arbitrary. To address this limitation, one approach is to measure the relative level of Shannon entropy compared to the maximum entropy distribution. Finally, emergence can be quantified by:
@@
where, $H_{max}$ is the maximized entropy of the system. If no prior information is available, the maximum entropy, denoted as $H_{max}$, corresponds to the Shannon entropy of an equal probability distribution. On the other hand, $H$ represents the Shannon entropy of the system at the final moment of a self-organization process. Additionally, $\Delta H_{view}$ represents the entropy increase during this process resulting from a change in the observer’s abstraction level. It should be noted that if the observer does not alter their abstraction level, this term would be zero. We can also normalize this quantity by dividing $H_{max}$ such that different features can be compared by the normalized quality. For a multivariate system, the authors suggest employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables. Then the authors apply this method to a simulated system of chickens. M. Tang and X. Mao applied this indicator on artificial society models  Tang and Mao (2014).
@@
Inspired by Moez Mnif and Christian Müller-Schloer’s work,  Fisch et al. (2010) suggests to use the divergence measure between two probability distributions to better quantify emergence. They understand emergence as an unexpected or unpredictable change of the distribution underlying the observed samples. However, this method has the defect of computational complexity and estimation accuracy. To conquer the problems,  Fisch et al. (2012) further proposed an approximating method by using Gaussian mixture model to estimate the density and introducing Mahalanobis distance to characterize the divergence between data and Gausian component, better results can be obtained. In Fisch et al. (2011), the authors systematically compare the three mentioned methods, and apply them on a simple test example. Another Shannon entropy based emergence measure is given by Holzer and de Meer et al. Holzer et al. (2008); Holzer and de Meer (2011). They consider a complex system as a self-organization process, in which different individuals interact with communications. Then emergence can be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
@@
Different from all the mentioned methods above, Teo et al. (2013); Szabo and Teo (2015) proposes a method to quantify emergence based on the idea of "the whole is greater than the sum of its parts", and it defines emergence from the interaction rules and states of agents instead of on the overall statistical measure of the whole system. Specifically, this measure consists of two terms that are subtracted from each other. The first term characterizes the collective states of the entire system, whereas the second term represents the summation of the individual states of all its constituent parts. This measure emphasize the emergence that arises from the interactions and collective behavior of the system. Then this method is tested on the example of bird flocks simulation.
@@ 3.5.1 Basic idea
The first quantitative emergence theory defined on Markov dynamics and causality measure based on intervention is Erik Hoel’s causal emergence theory.
In this framework, system properties can be characterized at various levels, ranging from micro to macro. If a system exhibits stronger causality at the macro level than at the micro level, it demonstrates causal emergence. Causality is reflected between successive states during the system’s evolution. The strength of a system’s causality reveals the extent to which its future state is influenced by its current state. The basic idea of causal emergence can be illustrated by Figure 5.
Figure 5: The basic idea of Erik Hoel’s causal emergence theory.
For example, statistical mechanics is a typical theory of causal emergence. On the micro-level, a huge number of molecules collide each other and exhibit random behaviors such that probabilistic language must be used to describe them. However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure and temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors. Therefore, thermodynamic laws have stronger causal effect than micro-level molecular dynamics.
Formal tools are used to describe the elements in Hoel’s causal emergence theory. Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro systems, and the inherent strength of causal effect of the Markov model can be measured with effective information (EI) which indicates how effectively a particular state influences the future state of a system.
EI is an intrinsic property of a system’s dynamics which can be quantified by the transitional probability matrix (TPM). A coarse-graining strategy is a function that can map a set of micro-states into a particular macro-state, and therefore a new dynamical model which can be described by TPM can naturally be derived from the micro-level TPM, and the effective information of the coarse-grained model can also be computed. The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase, and at a certain macroscopic scale of coarse-graining, the effective information reaches a maximum: this is the scale at which the system state has maximum causal power to predict future states in the most reliable and effective way.
To illustrate the causal emergence, let’s look at a particular Markov chain model with 4 possible states, whose transition probability matrix (TPM) is:
@@
In this model, if we take the first 3 states as a group 1 and the last state as group 2, anyone in group 1 can transit to all 3 states of the same group with equal probability at the next moment, while because there’s only 1 state in group 2, so the 4th state will always stay in its position. Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty is mainly from group 1.

### 25 Feb 2024 v3

@@ 3.3 G-Emergence Theory
G-emergence theory, proposed by Seth in 2008 Seth (2008), is one of the earliest works on a quantitative measure of emergence. His basic idea is to use nonlinear Granger causality to measure weak emergence in complex systems.
@@
Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$, then $B$ is said to be “Granger-cause” of $A$. This implies the existence of Granger causality between $A$ and $B$.
@@
When applying a bivariate autoregressive model for predictions, residual terms are included in the equations of the two variables. Then, the residuals can be utilized to quantify the extent of the causal effect in G-causality. The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals. One is the residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other is the residual of the full prediction model, as shown in Figure 3. In addition, in Seth (2008), the author also defined G-autonomous as the ability of past values in a time series to predict its own future values. And the degree of G-autonomous can be measured in a similar way as G-causality.
Figure 3: G-causality and G-autonomous.
@@
With these two basic notions in G-causality, as shown in Figure 4, the author defined G-emergence on macro-variables as follows: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if (i) $M$ is G-autonomous with respect to $\{m\}$, and (ii) $\{m\}$ is G-cause of $M$. The degree of G-emergence can also be quantitatively measured by multiplying the degree of G-autonomous of $M$ and the average of G-causes of $\{m\}$ on $M$.
Figure 4: G-autonomous.
@@
The author tested the G-emergence theory on the Boid model. This model is a famous artificial life model that simulates the flocking behavior of birds using three simple rules: cohesion, alignment, and separation Reynolds (1987). These basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled. The author found that by increasing the strength of cohesion, the G-emergence of the entire flock also increased. Here, the macro-variable is selected as the center of the flock (center of mass), and each bird is a micro-variable. The author also discovered a downward causality phenomenon in this simple model: the center of mass can be used to predict each individual bird. However, the author did not distinguish downward causation from other common causality in their work because Granger’s causality is not a real causal relationship.
@@
Seth’s G-emergence theory was the first attempt to quantify the emergence phenomenon via a causality measure. However, the causality measure that the author used was Granger causality, which is not a strict causal measure, and it also depends on the regression method to be used. Furthermore, the measure is defined on variables but not dynamics, which means the result depends on the selection of variables.
@@ 3.4 Other Quantitative Theories of Emergence
There are alternative means of measuring emergence that do not require relying on causality. Two different methods have been discussed by the scholars. One is to understand emergence as a process from disorder to order, and the other is to understand emergence from the perspective that “the whole is greater than the sum of parts”.
For example, Moez Mnif and Christian Müller-Schloer used Shannon entropy Mnif and Müller-Schloer (2011) to measure order and disorder. In a self-organized process, emergence occurs when there is an increase in order. This increase can be quantified by measuring the Shannon entropy difference between the initial state and the final state, denoted as $H_{start}-H_{end}$. However, this definition has two limitations. Firstly, the measurement of entropy, denoted as $H$, is dependent on the abstract level of observation. Therefore, it is necessary to account for the entropy increase resulting from a change in the observer’s abstract level. Secondly, the choice of the initial condition of the system is arbitrary. To address this limitation, one approach is to measure the relative level of Shannon entropy compared to the maximum entropy distribution. Finally, emergence can be quantified by:
@@
where $H_{max}$ is the maximized entropy of the system. If no prior information is available, the maximum entropy, denoted as $H_{max}$, corresponds to the Shannon entropy of an equal probability distribution. On the other hand, $H$ represents the Shannon entropy of the system at the final moment of a self-organization process. Additionally, $\Delta H_{view}$ represents the entropy increase during this process resulting from a change in the observer’s abstraction level. It should be noted that if the observer does not alter their abstraction level, this term would be zero. We can also normalize this quantity by dividing $H_{max}$ such that different features can be compared by the normalized quality. For a multivariate system, the authors suggested employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables. Then, the authors applied this method to a simulated system of chickens. M. Tang and X. Mao applied this indicator to artificial society models Tang and Mao (2014).
@@
Inspired by Moez Mnif and Christian Müller-Schloer’s work,  Fisch et al. (2010) suggested using the divergence measure between two probability distributions to better quantify emergence. They understood emergence as being an unexpected or unpredictable change of the distribution underlying the observed samples . However, this method suffers from computational complexity and estimation accuracy. To address these problems, Fisch et al. (2012) further proposed an approximating method using the Gaussian mixture model to estimate the density and introduced Mahalanobis distance to characterize the divergence between data and Gaussian components, leading to better results. In Fisch et al. (2011), the authors systematically compared the three aforementioned methods and applied them to a simple test example. Another Shannon entropy-based emergence measure was proposed by Holzer and de Meer et al. Holzer et al. (2008); Holzer and de Meer (2011). They considered a complex system as a self-organization process in which different individuals interact with communications. Emergence can then be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
@@
Unlike the aforementioned methods, Teo et al. (2013); Szabo and Teo (2015) proposed a method to quantify emergence based on the idea that “the whole is greater than the sum of its parts”, defining emergence from the interaction rules and states of agents instead of the overall statistical measure of the whole system. Specifically, this measure consists of two terms that are subtracted from each other. The first term characterizes the collective states of the entire system, whereas the second term represents the summation of the individual states of all its constituent parts. This measure emphasizes the emergence that arises from the interactions and collective behavior of the system. This method was then tested on the example of bird flock simulation.
@@ 3.5.1 Basic Idea
The first quantitative emergence theory based on Markov dynamics and causality measures through intervention was Erik Hoel’s causal emergence theory.
In this framework, system properties can be characterized at various levels, ranging from micro to macro. If a system exhibits stronger causality at the macro-level than at the micro-level, it demonstrates causal emergence. Causality is reflected between successive states during the system’s evolution. The strength of a system’s causality reveals the extent to which its future state is influenced by its current state. The basic idea of causal emergence is illustrated in Figure 5.
Figure 5: The basic idea of Erik Hoel’s causal emergence theory. The colored circles represent micro-states, whereas the colored squares represent macro-states.
For example, statistical mechanics is a typical theory of causal emergence. At the micro-level, a huge number of molecules collide and exhibit random behaviors such that probabilistic language must be used to describe them. However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure, temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors. Therefore, thermodynamic laws have a stronger causal effect than micro-level molecular dynamics.
Formal tools are used to describe the elements in Hoel’s causal emergence theory. Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro-systems. Additionally, the inherent strength of the causal effect of the Markov model can be measured with effective information (EI), indicating how effectively a particular state influences the future state of a system.
EI is an intrinsic property of a system’s dynamics and can be quantified by the transitional probability matrix (TPM). A coarse-graining strategy is a function that maps a set of micro-states into a particular macro-state, allowing for the derivation of a new dynamical model described by TPM from the micro-level TPM. The effective information of the coarse-grained model can also be computed. The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase. At a certain macroscopic scale of coarse-graining, the effective information reaches a maximum; this scale represents the point at which the system state has the maximum causal power to predict future states in the most reliable and effective way.
To illustrate causal emergence, let us look at a particular Markov chain model with four possible states, whose transition probability matrix (TPM) is:
@@
In this model, if we take the first three states as group 1 and the last state as group 2, anyone in group 1 can transition to any of the three states in the same group with equal probability at the next moment. However, because there is only one state in group 2, so the fourth state will always stay in its position. Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty mainly arises from group 1.